# Robust $Q$-Learning for finite ambiguity sets

### v2

@@ Abstract.
In this paper we propose a novel $Q$-learning algorithm allowing to solve distributionally robust Markov decision problems for which the ambiguity set of probability measures can be chosen arbitrarily as long as it comprises only a finite amount of measures. Therefore, our approach goes beyond the well-studied cases involving ambiguity sets of balls around some reference measure with the distance to reference measure being measured with respect to the Wasserstein distance or the Kullback–Leibler divergence. Hence, our approach allows the applicant to create ambiguity sets better tailored to her needs and to solve the associated robust Markov decision problem via a $Q$-learning algorithm whose convergence is guaranteed by our main result. Moreover, we showcase in several numerical experiments the tractability of our approach.
@@
Corresponding author, email:jul__\__ses@nus.edu.sg
@@ 1. Introduction
Markov decision problems are discrete time stochastic control problems involving an agent who can influence the environment by executing actions affecting the distribution of the realization of the subsequent state of the environment. In turn, depending on the state realization, the agent receives a feedback in form of a reward and decides on her next action. The objective of the agent is then to optimize the cumulative expected rewards obtained over an infinite time horizon by choosing the highest rewarding sequence of actions.
By construction, the described optimization problem depends crucially on the involved distributions for the state transitions. In practice, frequently in model-based approaches these distributions are estimated ([1], [35], [40]) from data or implicitly learned on-line in model-free approaches (see, e.g., [9], [42]). However chosen, the underlying model is highly prone to model misspecification, i.e., if the underlying distribution changes after training, the agent is trained to the wrong underlying distribution and successful execution of actions (i.e., receiving high rewards) typically will fail if the distribution change is too large. A solution to this problem is to take distributional uncertainty into account, i.e., to train the agent already by using an ambiguity set of probability distributions and train the agent according to the worst case measure contained in the ambiguity set, following the paradigm that during execution the distribution of the state transition might change but still is contained in a pre-defined ambiguity set. In this paper we follow this paradigm and propose a novel numerical algorithm to solve Markov decision problems while accounting for distributional uncertainty by considering arbitrary finite ambiguity sets.
@@ 1.1. Our Contribution
One frequently used numerical method to solve (non-robust) Markov decision problems that has turned out to work efficiently in practice is the so-called $Q$-learning algorithm which initially was introduced by Watkins in his PhD thesis, see [50], as well as [51] for the accompanying paper. For a general introduction to $Q$-learning algorithms we refer to [12] and [20], and for more example of associated applications, see among many others [3], [4], [10], [11], [17], [22], and [28].
Our contribution is to present a novel $Q$-learning algorithm that allows to take distributional uncertainty for the state transitions of Markov decision problems into account and which therefore respects the so called Knightian uncertainty ([21]), i.e., the uncertainty of having chosen the correct model.
We respect this model uncertainty by, instead of fixing a unique probability distribution for the state transitions, allowing the state transition between two subsequent states to be realized according to any distribution from a set of predetermined transition probabilities and then to consider the worst-case measure to compute rewards. Moreover, to account for the fact that transition probabilities may change over time, we consider a time-inhomogeneous formulation, i.e., we allow different transition probabilities in every instance of time as long as the transition probabilities are contained in the ambiguity set. To solve this robust formulation of a Markov decision process (as introduced in [31]) under the worst-case measure we develop a novel $Q$-learning algorithm.
The finite ambiguity set of transition probabilities can be designed by the applicant and its shape is, in particular, not restricted to measures that are in a certain sense close to a reference measure. This allows to consider different ambiguity sets as it were possible in the case of ambiguity sets that are defined as balls around reference measures. Put differently, using our approach, an applicant can simply include any possible scenarios in which she is interested and hence is able to control the shape and content of the ambiguity sets and hence the dynamics of the underlying Markov decision process that are deemed to be admissible. Moreover, in Section 6 we provide a proof guaranteeing convergence of our algorithm.
@@ 1.2. Related Work
While the study of distributionally robust optimization (DRO) problems and distributionally robust Markov decision problems and problems has been an active research topic already in the past decade (compare, e.g., [6], [7], [16], [24], [31], [32], [33], [37], [38], [44], [49], [52], [53], [54], and [55]), the discussion and construction of $Q$-learning algorithms to solve these sequential decision making problems numerically has only become an active research topic very recently.
In [7], [23], and [48] Q-learning algorithms for ambiguity sets that are constructed as balls around some reference measures with respect to the Kullback–Leibler divergence are discussed.
Moreover, in [29], the authors present a novel $Q$-learning algorithm for the specific case that the ambiguity set is given by a Wasserstein-ball around a reference measure.
To the best of our knowledge going beyond the aforementioned cases imposes a novelty, and has not been studied yet.
@@ 1.3. Structure
The remainder of this paper is as follows. In Section 2 we present the setting and the associated stochastic control problem. A numerical solution to this optimization problem via $Q$-learning is presented in Section 3 together with our main result guaranteeing convergence of the $Q$-learning algorithm. Section 4 discusses extensions of the presented setting to ambiguity sets with infinitely many probability measures and to continuous state spaces. In Section 5 we provide numerical experiments showcasing the tractability of our $Q$-learning algorithm. The proofs are reported in Section 6.
@@ 2.1. Setting
[…]
@@
$(x,a)$ $\twoheadrightarrow\mathcal{P}(x,a):=\left\{\mathbb{P}^{(1)}(x,a),\cdots,\mathbb{P}^{(N)}(x,a)\right\},\ \ \ \ \ \ N\in\mathbb{N},$
@@
$\mathbb{P}(B)=\sum_{x_{0}\in\mathcal{X}}...\sum_{x_{t}\in\mathcal{X}}...1\mathrm{l}_{B}((x_{t})_{t\in\mathbb{N}})...\mathbb{P}_{t-1}(x_{t-1};\{x_{t}\})...\mathbb{P}_{0}(x_{0};\{x_{1}\})\delta_{x}(\{x_{0}\}),\ \ \ \ \ \ B\in\mathcal{F}.$
@@ 2.2. Finite ambiguity sets and other ambiguity sets
The Q-learning approaches presented in [7], [23], and [29] provide solutions to account for a potential misspecification of a reference measure, by allowing deviations from it where the size of the deviations is measured by the respective distances (Kullback–Leibler distance and Wasserstein distance). However, these approaches allow not to control of which type the distributions in the ambiguity set are but simply consider all measures that are in a certain sense close to the reference measure. If an applicant is instead interested in controlling the distributions contained in the ambiguity set, for example, by allowing only for a specific type of parametric distributions with a finite set of possible parameters or by considering an asymmetric ambiguity set, the applicant can use the $Q$-learning approach presented in this paper. Another relevant situation is the consideration of multiple estimated models / probability measures obtained for exampled via different estimation methods. To account for all of these estimations, an applicant can simply construct an ambiguity set containing all estimated probability measures.
@@
The main difference therefore is that an applicant of the approach presented in this paper can control exactly what types of distributions are deemed possible whereas the approaches from [7], [23], and [29] account for a misspecification of one single reference measure.
@@ 2.3. Specification of the optimization problem
After execution of an action $a_{t}(X_{t})$, the agent receives a feedback on the quality of the chosen action in terms of a reward $r(X_{t},a_{t}(X_{t}),X_{t+1})$ where $r$ denotes some reward function $r:\mathcal{X}\times A\times\mathcal{X}\rightarrow\mathbb{R}$. The robust optimization problem consists then, for every initial state $x\in\mathcal{X}$, in maximizing the expected value of $\sum_{t=0}^{\infty}\alpha^{t}r(X_{t},a_{t}(X_{t}),X_{t+1})$ under the worst case measure from $\mathfrak{P}_{x,\textbf{a}}$ over all possible policies $\textbf{a}\in\mathcal{A}$, where $\alpha\in\mathbb{R}$ is some discount factor accounting for time preferences of rewards. Therefore, the value function
@@ 2.4. Dynamic Programming
To solve (2.3) directly is typically a non-feasible optimization problem as it means to find directly (infinite-dimensional) solutions over the whole time horizon. However, due to its time-homogeneous structure, solving the infinite time horizon problem can be reduced to a one time step problem which turns out to be tractable. To see this, we consider the one step optimization problem
(2.4) $\mathcal{T}V(x):=\max_{a\in A}\min_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right],\ \ \ \ x\in\mathcal{X}$
where $\mathcal{X}\ni x\mapsto V(x)$ is the value function defined in (2.3). Analogue to the definition of the non-robust $Q$-value function, see e.g. [50], we define the robust optimal Q-value function by
@@
allowing us to interpret $Q^{*}(x,a)$ as the quality of executing action $a$ when in state $x$ (under the worst case measure) with the best possible action leading to $\mathcal{T}V(x)$. By using the above definitions we then obtain the dynamic programming equation $\mathcal{T}V=V$ explaining the notion of quality. This fixed-point equation is also famously known in its non-robust formulation as Bellman equation, compare e.g. [13, p.164] for a non-robust version.
@@
Assume $10<\alpha<$ 1. Then, for all $x\in\mathcal{X}$ we have
@@ 3. Robust Q-Learning Algorithm
In this section we propose with Algorithm 1 a novel distributionally robust $Q$-learning algorithm and provide in Theorem 3.1 conditions for its convergence.
@@
  Input:
   State space $\mathcal{X}\subseteq\mathbb{R}^{d}$;  Initial state $x_{0}\in\mathcal{X}$;  Action space $A\subseteq\mathbb{R}^{m}$;
   Reward function $r$;  Discount factor $\alpha\in(0,1)$;
   Ambiguity set $\mathcal{X}\times A\ni(x,a)\twoheadrightarrow\mathcal{P}(x,a):=\left\{\mathbb{P}^{(1)}(x,a),\ldots,\mathbb{P}^{(N)}(x,a)\right\}$;
   Policy $(a_{t})_{t}\in\mathcal{A}$;  Sequence of learning rates $(\widetilde{\gamma_{t}})_{t\in\mathbb{N}}\subseteq[0,1]$;
@@
(3.1) $k_{t}^{*}\in\operatorname*{arg\,min}_{k=1,\ldots,N}\mathbb{E}_{\mathbb{P}^{(k)}(X_{t},a_{t}(X_{t}))}\left[r(X_{t},a_{t}(X_{t}),X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right];$
     Sample $X_{t+1}\sim\mathbb{P}^{(k_{t}^{*})}\left(X_{t},a_{t}(X_{t})\right)$;
@@
(3.2) $Q_{t+1}(x,a)\leftarrow\begin{cases}(1-\widetilde{\gamma}_{t})\cdot Q_{t}(x,a)+\widetilde{\gamma}_{t}\cdot\left(r(x,a,X_{t+1})+\alpha\max\limits_{b\in A}Q_{t}(X_{t+1},b)\right)&\text{if }(x,a)=\left(X_{t},a_{t}(X_{t})\right),\\ Q_{t}(x,a);&\text{else}\end{cases}$
@@
Given a policy $\textbf{a}\in\mathcal{A}$ and some initial state $x_{0}\in\mathcal{X}$ we define a proability measure by
(3.3) $\mathbb{P}_{x_{0},\textbf{a}}:=\delta_{x_{0}}\otimes\mathbb{P}_{0}^{(k_{0}^{*})}(\cdot,a_{0}(\cdot))\otimes\mathbb{P}_{1}^{(k_{1}^{*})}(\cdot,a_{1}(\cdot))\otimes\cdots\in\mathfrak{P}_{x_{0},\textbf{a}},$
where $k_{t}^{*}$ is as defined in (3.1). Note that the probability measure defined in  (3.3) is the probability measure according to which the states are sampled when applying Algorithm 1.
Our following main result now shows that the function $(Q_{t})_{t\in\mathbb{N}}$ obtained as the output of Algorithm 1 converges pointwise almost surely towards the optimal robust $Q$-value function $Q^{*}$ defined in (2.5).
@@
(3.4) $\gamma_{t}(x,a,X_{t}):=1\mathrm{l}_{\left\{(x,a)=\left(X_{t},~{}a_{t}(X_{t})\right)\right\}}\cdot\widetilde{\gamma}_{t}.$
Moreover, let $10<\alpha<$ 1 and let $(x_{0},\textbf{a})\in\mathcal{X}\times\mathcal{A}$ such that
@@
[…]
@@
    Note that in many situations one can improve the convergence speed by replacing $\widetilde{\gamma}_{t}$ with $\widetilde{\gamma}_{\operatorname{visits}_{t}(x,a)}$, where we initially set $\operatorname{visits}_{0}\equiv 0$, and then for all $(x,a)\in\mathcal{X}\times A$ and $t=1,2,\dots,$
    $\operatorname{visits}_{t}(x,a)\leftarrow\left\{\begin{array}[]{ll}\operatorname{visits}_{t}(x,a)+1&\text{if}\ (x,a)=\left(X_{t},~{}a_{t}(X_{t})\right),\\ \operatorname{visits}_{t}(x,a)&\text{else},\end{array}\right.$
    compare, e.g., [25]. The idea behind this modification is to account for the exploration-exploitation trade off by reducing the learning rate the more often we have visited a state and hence the more confident the algorithm is w.r.t.  choosing the correct optimal action, and conversely to apply a larger learning rate if the state-action pair was visited less often until time $t$.
@@
    $\mathcal{X}\ni x\mapsto a_{t}(x):=\begin{cases}\operatorname{argmax}_{b\in B}Q_{t}(x,b)&\text{with probability }1-\epsilon_{\operatorname{greedy}},\\ a\sim\mathcal{U}(A)&\text{with probability }\epsilon_{\operatorname{greedy}},\end{cases}$

### v3

# Q-Learning under Finite Model Uncertainty

@@ Abstract.
We propose a robust Q-learning algorithm for Markov decision processes under model uncertainty when each state–action pair is associated with a finite ambiguity set of candidate transition kernels. This finite-measure framework enables highly flexible, user-designed uncertainty models and goes beyond the common KL/Wasserstein-ball formulations. We establish almost sure convergence of the learned Q-function to the robust optimum, and derive non-asymptotic high-probability error bounds that separate stochastic approximation error from transition-kernel estimation error. Finally, we show that Wasserstein-ball and parametric ambiguity sets can be approximated by finite ambiguity sets, allowing our algorithm to be used as a generic solver beyond the finite setting.
@@
Corresponding author, email:jul_\_ses@nus.edu.sg
@@ 1. Introduction
Markov decision processes (MDPs) are a fundamental framework for sequential decision making under uncertainty, and, in particular, for reinforcement learning. An agent repeatedly observes the state of a stochastic environment, selects an action, and receives a reward, with the goal of maximizing cumulative discounted rewards over an infinite time horizon. Central to this formulation is the specification of the state transition probabilities, which determine how actions influence the evolution of the system.
In practice, these transition probabilities are rarely known and must be estimated from data ([1], [44], [50]) or implicitly learned through interaction with the environment (see, e.g., [10], [53]). As a consequence, classical (non-robust) reinforcement learning methods are highly sensitive to model misspecification: policies trained under an incorrect or outdated transition model may perform poorly when deployed, even if the deviation from the training model is moderate. This issue is particularly acute in non-stationary environments or in applications where data are scarce and multiple competing models are plausible. A natural way to address this challenge is to incorporate model uncertainty directly into the training phase. In distributionally robust Markov decision problems ([38]), the transition kernel is assumed to belong to an ambiguity set of probability measures, and the agent optimizes performance under the worst-case model within this set. This paradigm reflects Knightian uncertainty ([26]) and provides policies that are robust against adverse but plausible deviations of the true dynamics.
Most existing robust reinforcement learning approaches specify ambiguity sets as metric balls, such as Kullback–Leibler ([7], [29]) or Wasserstein balls ([39]), centered at a reference transition kernel. While mathematically convenient, this construction implicitly enforces a neighborhood-based and typically symmetric notion of uncertainty. As a result, it does not allow the decision maker to precisely control which alternative transition models are considered plausible, nor to express asymmetric or scenario-based uncertainty.
In many applications, however, uncertainty is better described by a finite collection of competing models. These may arise from different estimation procedures, stress scenarios, regime-specific dynamics, or expert judgments. In such settings, it is often desirable to protect explicitly against a small number of adverse scenarios rather than against all distributions within a metric neighborhood of a reference model. This motivates the study of robust control problems with finite, user-specified ambiguity sets.
To illustrate the systematic difference, consider an agent repeatedly betting on the outcome of a coin toss game, where the state records recent outcomes and the agent chooses whether to bet on an increase or decrease of the next realization. Suppose the true data-generating process is unknown but believed to belong to a small set of plausible models, such as a fair or a moderately biased coin. A non-robust agent trained under a single reference model may perform well when this model is correct, but can suffer systematic losses under alternatives. Wasserstein-robust approaches mitigate this by protecting against all distributions in a neighborhood of the reference model; however, this typically introduces excessive conservatism by including many unintended intermediate distributions. In contrast, finite ambiguity sets allow the decision maker to specify explicitly the competing scenarios and to optimize against their worst case. As shown in Section 6, this scenario-based robustness yields qualitatively different policies and can outperform both non-robust and Wasserstein-robust strategies when uncertainty is concentrated on a small number of concrete models.
In this paper, we develop a robust $Q$-learning framework for Markov decision problems in which each state–action pair is associated with a finite set of candidate transition kernels. We allow the transition kernel selected by nature to vary over time, as long as it remains within the prescribed ambiguity set, leading to a time-inhomogeneous worst-case formulation. This setting preserves the dynamic programming principle while enabling highly flexible modeling of uncertainty.
@@ 1.1. Our contribution
Our main contributions can be summarized as follows.
  1. 1.)
    We introduce a distributionally robust MDP framework based on finite ambiguity sets of transition probabilities. This approach enables scenario-based robustness and allows the decision maker to specify explicitly which alternative models are deemed plausible, without being restricted to metric balls around a reference measure.
  2. 2.)
    We propose a novel robust $Q$-learning algorithm for this setting. The algorithm requires only generative access to each candidate transition kernel and does not assume explicit knowledge of the underlying distributions. At each iteration, the learner evaluates actions under the empirically worst-case model and updates the Q-function accordingly.
  3. 3.)
    We provide a rigorous theoretical analysis of the proposed algorithm. We establish almost sure convergence of the learned Q-function to the robust optimal Q-function and derive non-asymptotic high-probability error bounds. These bounds naturally decompose into a stochastic approximation error and an additional term reflecting the statistical error arising from estimating the candidate transition kernels. In the special case where the kernels are known, this latter term vanishes.
  4. 4.)
    We show that a broad class of ambiguity sets, including Wasserstein balls and parametric families of transition kernels, can be approximated by sequences of finite ambiguity sets. These results allow the proposed algorithm to serve as a generic numerical solver beyond the finite setting.
  5. 5.)
    We show tractability of our approach through two examples: one coin toss example allowing to fully understand the presented approach and how it distinguishes from other related approaches as well as a stock allocation example using real data.
@@ 1.2. Related work
Distributionally robust optimization and robust Markov decision problems have been studied extensively over the past decade (compare, e.g., [7], [8], [17], [31], [38], [41], [42], [47], [48], [55], [60], [64], [65], [66], and [67]). This literature has primarily focused on establishing dynamic programming principles under model uncertainty, characterizing robust value functions, and developing tractable solution methods for specific classes of ambiguity sets. A central modelling choice in these works is the specification of uncertainty via ambiguity sets of transition kernels, typically designed to capture deviations from a reference model in a principled manner.
Classical Q-learning for Markov decision problems was introduced by Watkins ([62] [61]), and has since become one of the cornerstone algorithms in reinforcement learning. Comprehensive introductions can be found in [13] and [24], and a wide range of applications has been studied, including control, finance, and operations research (see, e.g., [3], [4], [11], [12], [21], [27], and [37]). These approaches, however, assume that the underlying transition kernel is fixed and correctly specified, making them sensitive to model misspecification.
Only more recently has attention shifted to robust Q-learning methods that explicitly account for uncertainty in the transition dynamics. Existing approaches predominantly consider ambiguity sets defined as metric balls—most notably Kullback–Leibler balls ([8], [29], and [59]) and Wasserstein balls ([38])—around a reference transition kernel. These formulations exploit convex duality and the structure of the underlying metric to obtain tractable algorithms and theoretical guarantees. At the same time, metric-ball ambiguity sets impose a neighbourhood-based and typically symmetric notion of uncertainty, which implicitly includes many transition models that may not be considered plausible by a decision maker.
In contrast, many practical applications call for scenario-based uncertainty modelling, where ambiguity is naturally described by a finite collection of competing transition models arising from different estimators, stress scenarios, regimes, or expert assessments. Finite ambiguity sets are also particularly natural in financial applications, where uncertainty is often driven by a small number of stress scenarios or regime-dependent models rather than diffuse perturbations of a reference measure, see, e.g. [20], [25], and [45].
To the best of our knowledge, robust $Q$-learning methods that allow for arbitrary finite collections of candidate transition kernels, without relying on metric-ball constructions or convexity assumptions, have not been developed. The present work fills this gap by providing an algorithmic and theoretical treatment of robust $Q$-learning under finite model uncertainty, thereby enabling a flexible, user-designed approach to robustness that complements existing metric-based methods.
@@ 1.3. Structure
The remainder of the paper is organized as follows. Section 2 introduces the robust Markov decision problem with finite ambiguity sets and establishes the associated dynamic programming formulation. In Section 3, we present the robust $Q$-learning algorithm and prove its almost sure convergence. Section 4 shows how ambiguity sets consisting of infinitely many probability measures can be approximated by finite sets, including Wasserstein and parametric ambiguity models. Section 5 outlines an extension to continuous state spaces using function approximation and deep $Q$-learning. Section 6 provides numerical experiments illustrating the tractability and qualitative behavior of the proposed approach. All proofs are collected in Section 7.
@@ 2.1. Setting
In contrast to classical non-robust Markov decision problems, we work under the paradigm that the precise state transition probability is not known but instead, to account for model uncertainty, contained in an ambiguity set of finitely many transition probabilities. In the following, let $\mathcal{M}_{1}(\mathcal{X})$ denote the set of probability measures on $(\mathcal{X},\mathcal{F})$, and let $\tau_{0}$ denote the topology of weak convergence [^1] For any $\mu\in\mathcal{M}_{1}(X)$, $(\mu_{n})_{n\in\mathbb{N}}\subseteq\mathcal{M}_{1}(X)$ and $C_{b}(X,\mathbb{R})$ denoting the space of continuous and bounded functions mapping from the space $X$ to $\mathbb{R}$, we have: $\underset{n\rightarrow\infty}{\mu_{n}\longrightarrow^{\tau_{0}}\mu}\Leftrightarrow\lim\limits_{n\to\infty}\int gd\mu_{n}=\int gd\mu,\ \text{ for all }g\in C_{b}(X,\mathbb{R}).$ . An ambiguity set of $N\in\mathbb{N}$ probability measures in dependence of a state-action pair is then modeled by a set-valued map given by
@@
$(x,a)$ $\twoheadrightarrow\mathcal{P}(x,a)=\left\{\mathbb{P}^{(1)}(x,a),\cdots,\mathbb{P}^{(N)}(x,a)\right\},$
@@
$\mathbb{P}(B)=\sum_{x_{0}\in\mathcal{X}}...\sum_{x_{t}\in\mathcal{X}}...1\hskip-3.0pt\mathrm{l}_{B}((x_{t})_{t\in\mathbb{N}})...\mathbb{P}_{t-1}(x_{t-1};\{x_{t}\})...\mathbb{P}_{0}(x_{0};\{x_{1}\})\delta_{x}(\{x_{0}\}),\qquad B\in\mathcal{F}.$
@@ 2.2. Finite ambiguity sets and other ambiguity sets
The $Q$-learning approaches presented in [8], [29], and [39] provide solutions to account for a potential misspecification of a reference measure, by allowing deviations from it where the size of the deviations is measured by the respective distances (Kullback–Leibler distance and Wasserstein distance). However, these approaches allow not to control of which type the distributions in the ambiguity set are but instead consider all measures that are in a certain sense close to the reference measure. If a decision maker is instead interested in controlling the distributions contained in the ambiguity set, for example, by allowing only for a specific type of parametric distributions with a finite set of possible parameters or by considering an asymmetric ambiguity set, the decision maker can use the $Q$-learning approach presented in this paper. Another relevant situation is the consideration of multiple estimated models / probability measures obtained for exampled via different estimation methods. Wishing to take into account all of these estimations, a decision maker can construct an ambiguity set containing all estimated probability measures.
@@
The main difference therefore is that a decision maker using the approach presented in this paper can precisely control what types of distributions are deemed possible whereas the approaches from [8], [29], and [39] account for a misspecification of one single reference measure.
@@ 2.3. Specification of the optimization problem
After execution of an action $a_{t}(X_{t})$, the agent receives a feedback on the quality of the chosen action in terms of a reward $r(X_{t},a_{t}(X_{t}),X_{t+1})$ where $r$ denotes some reward function $r:\mathcal{X}\times A\times\mathcal{X}\rightarrow\mathbb{R}$. The robust optimization problem consists then, for every initial state $x\in\mathcal{X}$, in maximizing the expected value of $\sum_{t=0}^{\infty}\alpha^{t}r(X_{t},a_{t}(X_{t}),X_{t+1})$ under the worst case measure from $\mathfrak{P}_{x,\textbf{a}}$ over all possible policies $\textbf{a}\in\mathcal{A}$, where $\alpha\in(0,1)$ is some discount factor accounting for time preferences of rewards. Therefore, the value function
@@ 2.4. Dynamic programming
To solve (2.3) directly is typically a non-feasible optimization problem as it means to find directly (infinite-dimensional) solutions over the whole time horizon. However, due to its time-homogeneous structure, solving the infinite time horizon problem can be simplified significantly to a one time step problem which turns out to be tractable. To this end, we consider the one time step optimization problem
(2.4) $\mathcal{X}\ni x\mapsto\mathcal{T}V(x):=\max_{a\in A}\min_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right],$
where $\mathcal{X}\ni x\mapsto V(x)$ is the value function defined in (2.3). By using the above definitions one can show (see Proposition 2.1) that the dynamic programming equation $\mathcal{T}V=V$ holds, implying that solving the fixed point equation $\mathcal{T}V=V$ is sufficient to compute the optimal value function. This fixed-point equation is also famously known in its non-robust formulation as Bellman equation, compare, e.g. [14, p.164] for a non-robust version.
To determine optimal actions, analogous to the definition of the non-robust $Q$-value function (see e.g. [62]), we further define the robust optimal $Q$-value function by
@@
allowing us to interpret $Q^{*}(x,a)$ as the quality of executing action $a$ when in state $x$ (under the worst case measure) with the best possible action leading to $V(x)$ explaining the notion of quality.
@@
Assume $0<\alpha<1$. Then, for all $x\in\mathcal{X}$ we have
@@
Using Proposition 2.1, the goal of our $Q$-learning algorithm presented in the next section will be to algorithmically determine $Q^{*}$ which in turn allows to infer optimal actions in every given state.
@@ 3. Robust Q-Learning Algorithm
In this section, we introduce Algorithm 1, a novel distributionally robust $Q$-learning algorithm. We establish sufficient conditions for its convergence in Theorem 3.1 and derive explicit convergence rates in Subsection 3.2.
3.1. The robust Q-Learning algorithm and its convergence
We do not assume explicit knowledge of the candidate transition kernels in $\mathcal{P}$. Instead, we only require generative access: for every $(x,a)\in\mathcal{X}\times A$ and each $k\in\{1,\dots,N\}$, the learner can generate independent samples from the transition kernel $\mathbb{P}^{(k)}(x,a)$. Thus, the algorithm only requires access to $N$ independent simulators, a standard assumption in generative-model reinforcement learning.
@@ 3.1. The robust Q-Learning algorithm and its convergence
Input:  State space $\mathcal{X}\subseteq\mathbb{R}^{d}$;  Initial state $x_{0}\in\mathcal{X}$;  Action space $A\subseteq\mathbb{R}^{m}$;  Reward function $r$;  Discount factor $\alpha\in(0,1)$;  Ambiguity set $\mathcal{P}(x,a)=\left\{\mathbb{P}^{(1)}(x,a),\ldots,\mathbb{P}^{(N)}(x,a)\right\}$;  Policy $(a_{t})_{t}\in\mathcal{A}$;  Sequence of learning rates $(\widetilde{\gamma_{t}})_{t\in\mathbb{N}}\subseteq[0,1]$;
@@
 Initialize $\mathbb{P}_{0}^{(k)}(x,a)(y)=\frac{1}{|\mathcal{X}|}$ for all $(x,a,y)\in\mathcal{X}\times A\times\mathcal{X}$;
 Initialize $N_{0}(x,a)=0$ for all $(x,a)\in\mathcal{X}\times A$;
@@
(3.1) $k_{t}^{*}\in\operatorname*{arg\,min}_{k=1,\ldots,N}\left\{\sum_{y\in\mathcal{X}}\left(r(X_{t},a_{t}(X_{t}),y)+\alpha\max_{b\in A}Q_{t}(y,b)\right)\mathbb{P}_{t}^{(k)}(X_{t},a_{t}(X_{t}))(y)\right\};$
  Sample $X_{t+1}^{(k)}\sim\mathbb{P}^{(k)}\left(X_{t},a_{t}(X_{t})\right)$ independently for $k=1,\dots,N$;
  Set $X_{t+1}:=X_{t+1}^{(k_{t}^{*})}$;
  Set
(3.2) $\mathbb{P}_{t+1}^{(k)}(x,a)(y):=\frac{\sum_{s=0}^{t}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}),X_{s+1}^{(k)})=(x,a,y)\}}}{\sum_{s=0}^{t}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}}\text{ for all }k\in\{1,\dots,N\},~~(x,a,y)\in\mathcal{X}\times A\times\mathcal{X};$
@@
(3.3) $Q_{t+1}(x,a)\leftarrow\begin{cases}(1-\widetilde{\gamma}_{N_{t}(x,a)})\cdot Q_{t}(x,a)+\widetilde{\gamma}_{N_{t}(x,a)}\cdot\left(r(x,a,X_{t+1})+\alpha\max\limits_{b\in A}Q_{t}(X_{t+1},b)\right)&\text{if }(x,a)=\left(X_{t},a_{t}(X_{t})\right),\\ Q_{t}(x,a);&\text{else};\end{cases}$
and update the number of visit times $N_{t}$ via
$N_{t+1}(x,a)\leftarrow\left\{\begin{array}[]{ll}N_{t}(x,a)+1&\text{if}\ (x,a)=\left(X_{t},~a_{t}(X_{t})\right),\\ N_{t}(x,a)&\text{else};\end{array}\right$.
@@
Given a policy $\textbf{a}\in\mathcal{A}$ and some initial state $x_{0}\in\mathcal{X}$ we further define a probability measure by
(3.4) $\mathbb{P}_{x_{0},\textbf{a}}:=\delta_{x_{0}}\otimes\mathbb{P}^{(k_{0}^{*})}(\cdot,a_{0}(\cdot))\otimes\mathbb{P}^{(k_{1}^{*})}(\cdot,a_{1}(\cdot))\otimes\cdots\in\mathfrak{P}_{x_{0},\textbf{a}},$
where $k_{t}^{*}$ is defined in (3.1) and denotes the index of the transition kernel that attains the worst-case expected continuation value (based on the empirical estimate) at time $t$. The measure $\mathbb{P}_{x_{0},\mathbf{a}}$ thus characterizes the (unknown) law of the state process generated by Algorithm 1 under policy $\mathbf{a}$.
Our main result establishes that the sequence $(Q_{t})_{t\in\mathbb{N}}$ produced by Algorithm 1 converges pointwise (in $(x,a)$), $\mathbb{P}_{x_{0},\mathbf{a}}$-almost surely, to the optimal robust $Q$-value function $Q^{*}$ defined in (2.5).
@@
(3.5) $\gamma_{t}(x,a,X_{t}):=1\hskip-3.0pt\mathrm{l}_{\left\{(x,a)=\left(X_{t},~a_{t}(X_{t})\right)\right\}}\cdot\widetilde{\gamma}_{N_{t}(x,a)}.$
where
$N_{t}(x,a):=\sum_{s=0}^{t}\mathbf{1}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}$
is the number of visits to $(x,a)\in\mathcal{X}\times A$ up to time $t$. Moreover, let $0<\alpha<1$ and let $(x_{0},\textbf{a})\in\mathcal{X}\times\mathcal{A}$ such that
@@
Then, we have  $\text{ for all }(x,a)\in\mathcal{X}\times A$ that
@@
Remark 3.2.
If one has explicit knowledge of the distribution $\mathbb{P}^{(k)}$ (and not only a ready-to-use simulator that can sample from the distribution) for each candidate model $k=1,\dots,N$, then (3.1) can be replaced by
(3.8) $k_{t}^{*}\in\operatorname*{arg\,min}_{k=1,\ldots,N}\left\{\sum_{y\in\mathcal{X}}\left(r(X_{t},a_{t}(X_{t}),y)+\alpha\max_{b\in A}Q_{t}(y,b)\right)\mathbb{P}{(k)}(X_{t},a_{t}(X_{t}))(y)\right\};$
and there is no need to learn $\mathbb{P}_{t}$ via (3.2).
@@
    The idea behind replacing $\widetilde{\gamma}_{t}$ with $\widetilde{\gamma}_{N_{t}(x,a)}$ is to account for the exploration-exploitation trade off by reducing the learning rate the more often we have visited a state and hence the more confident the algorithm is w.r.t.  choosing the correct optimal action, and conversely to apply a larger learning rate if the state-action pair was visited less often until time $t$. This effectively enhances the convergence speed, compare, e.g., [32].
@@
    $\mathcal{X}\ni x\mapsto a_{t}(x):=\begin{cases}\operatorname{argmax}_{b\in A}Q_{t}(x,b)&\text{with probability }1-\epsilon_{\operatorname{greedy}},\\ a\sim\mathcal{U}(A)&\text{with probability }\epsilon_{\operatorname{greedy}},\end{cases}$